OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = -3/8 and passing through (5, -4)

OpenStudy (anonymous):

formula for straight line: y-y1 = m(x-x1) sub coordinates with x1 and y1

OpenStudy (anonymous):

I need the answer!

OpenStudy (anonymous):

im not sure what the standard form is sorry, but the equation would be y+4=-3/8(x-5)